During an effusion experiment, it took 75 seconds for a certain number of moles of an unknown gas to pass through a tiny hole. Under the same conditions, the same number of moles of oxygen gas passed through the hole in 30 seconds. What is the molar mass of the unknown gas?


1
Expert's answer
2021-10-30T11:15:56-0400

"t_u = 75 \;s \\

t_{O_2} = 30 \;s \\

M(O_2) = 32.0 \;g/mol \\

\sqrt{\frac{M_2}{M_1}} = \frac{t_2}{t_1} \\

\sqrt{\frac{32.0}{M_u}} = \frac{30.0}{75.0} \\

\sqrt{\frac{32.0}{M_u}} = 0.4 \\

\frac{32.0}{M_u} = 0.4^2 = 0.16 \\

M_u = \frac{32.0}{0.16} = 200 \;g/mol"

Answer: 200 g/mol